# Interpretability in Parameter Space: Minimizing Mechanistic Description Length with Attribution-based Parameter Decomposition

### v3

@@ 1 Introduction
To this end, we introduce Attribution-based Parameter Decomposition (APD), a method that decomposes neural network parameters into components that are optimized for these three properties. In brief, APD involves decomposing the parameter vector of any neural network into a sum of parameter components. They are optimized such that they sum to the target parameters while only a minimal number of them are necessary for the causal process that computes the network’s output for any given input. They are also optimized to be less complex individually than the entire network, in that they span as few directions in activation space as possible across all layers.
@@ 2.2 Identifying networks’ mechanisms using Attribution-based Parameter Decomposition
APD aims to minimize the total description length of the mechanistic components used by the network per data point over the training dataset. It decomposes the network’s parameters $\theta^{*}\in\mathbb{R}^{N}$ into a set of parameter components and directly optimizes them to be faithful, minimal, and simple. A more detailed discussion of how our method relates to the Minimum Description Length principle can be found in Appendix A.2.
@@
    Attribution step: We want estimate the causal importance of each parameter component $P_{c}$ for the network’s output on each datapoint $f_{\theta^{*}}(x)$. In this step, we therefore calculate the attributions of each parameter component with respect to the outputs, $A_{c}(x)\in\mathbb{R}$. It would be infeasibly expensive to compute this exactly, since it would involve a large number of causal interventions, requiring one forward pass for every possible combination of component ablations. We therefore use an approximation. In this work, we use gradient attributions (Mozer and Smolensky, 1988; Molchanov et al., 2017; Nanda, 2022a; Syed et al., 2023), but other attribution methods may work as well. This step therefore involves one forward pass with the target model to calculate the output and one backward pass to compute the attributions with respect to the parameters, which are used to calculate the attributions with respect to each parameter component (Equation 8).
@@ 3.2 Toy Model of Compressed Computation
The target network is a residual MLP, consisting of a residual stream width of $d_{\rm resid}=1000$, a single MLP layer of width $d_{\rm mlp}=50$, a fixed, random embedding matrix with unit norm rows $W_{E}$, an unembedding matrix $W_{U}=W_{E}^{\top}$, and $100$ input features. See Figure 5 for an illustration of the network architecture.
@@
The large residual stream $d_{\rm resid}=1000$ was chosen as the trained target network performed better than the naive monosemantic baseline in this setting (small values of $d_{\rm resid}$ lead to higher interference and thus worse model performance). We chose fixed, instead of trained, embedding matrices to make it simpler to calculate the optimal monosemantic baseline and to simplify training.
@@ 4.1 Addressing issues that seem challenging from an activations-first perspective
    ‘Features as arbitrary functions’, but this fails to distinguish between features that appear to be fundamental abstractions (e.g. a ‘cat feature’) and those that don’t (e.g. a ‘cat+car’ feature).
@@
In our work, we decompose neural networks into parameter components that minimize mechanistic description length, which we define as the network’s ‘mechanisms’. A network’s mechanisms are not equivalent to its ‘features’. However, defining a network’s features as ‘properties of the input that activate particular mechanisms’ seems to overcome the definitional issues above. In particular, it overcomes the issues in Definition 1 because we expect mechanisms, which are simple, minimally active components, to prefer to learn ‘cat components’ rather than ‘cat+car components’, since the latter is neither simpler (t requires machinery that can detect both cats and cars) nor minimal ( cat+car detector will activate more often than either a cat or car detector). Additionally, it does not rely on a notion of human interpretability, thus overcoming the issue with Definition 2. It also seems to overcome the issues of Definition 3, since the definition is not circular and should also be able to identify multidimensional mechanisms (and hence multidimensional features that activate them), although we leave this for future work.
@@
Neural network representations appear not to neatly map to individual architectural components such as individual neurons, attention heads, or layers. Representations often appear to be spread across various architectural components, as in attention head superposition (Jermyn et al., 2023) or cross layer distributed representations (Yun et al., 2021; Lindsay et al., 2024). It is unclear how best to adapt SDL to each of these settings in order to tell concise mechanistic stories (Mathwin et al., 2024; Wynroe and Sharkey, 2024; Lindsay et al., 2024). A more general approach that requires no adaptation would be preferred. Interpretability in parameter suggests a way to overcome this problem in general, since any architecture can in theory be decomposed into directions in parameter space without the method requiring adaptation.
@@ 5 Conclusion
This work introduces Attribution-based Parameter Decomposition (APD) as a fundamental shift in mechanistic interpretability: instead of analysing neural networks through their activations, we demonstrate that directly decomposing parameter space can reveal interpretable mechanisms that are faithful, minimal, and simple. Our approach suggests solutions to long-standing problems in mechanistic interpretability, including identifying minimal circuits in superposition, providing a conceptual foundation for ’features’, enabling better understanding of ’feature geometry’, and serving as an architecture-agnostic approach to neural network decomposition.
@@ Appendix A More detailed description of the APD method
A.1 Component attributions
We want most parameter components $P_{c}\in\mathbb{R}^{N}$ to not be ‘used’ on any one network input $x$, in the sense that we can ablate all but a few of them without changing the outputs of the network.
In other words, if we initialize a new network with a parameter vector $\kappa(x)$ composed of only a few of the most ‘used’ parameter components, we should have:
@@ A.1 Component attributions
Since the ‘inactive’ components are supposed to play no meaningful role in the computation of the output, we should also be able to ablate them (including partial ablation - see the stricter definition of ’inactive’ below), and still get the same result.
Directly checking that this condition is satisfied would be computationally intractable. Instead, we try to estimate whether the condition is approximately satisfied by calculating attributions of the output to each component. Currently, we do this using gradient attributions (Finlayson et al., 2021; Molchanov et al., 2017; Nanda, 2022b). This estimates the effect of ablating $P_{c}$ as:
@@
A stricter definition of ‘inactive’ parameter components
In general, we should get the same network output for any parameter configuration everywhere along any monotonic ‘ablation curve’ $\gamma_{c}(x,t)$ where $t\in[0,1]$:
$f(x|\kappa(x,t))\approx f(x|\theta^{*})\,.$
This is a stricter definition of ‘inactive’ that seeks to exclude cases like components $P_{c}$ and $P_{c^{\prime}}$, both being ‘active’ and important but canceling each other out. Without this stricter condition, we could have pathological solutions to the optimisation problem. For example, if $P^{\prime}_{1},\dots,\,P^{\prime}_{C}$ are a large set of parameter vectors for expert networks specialized to particular tasks the target network is capable of, completely unrelated to the target network $\theta^{*}$, we could set
$P_{c}=P^{\prime}_{c}+\frac{1}{C}\left(\theta^{*}-\sum_{c^{\prime}}P^{\prime}_{c^{\prime}}\right)\,.$ (10)
@@
Equations 5 and 9 together define what we mean when we say that we want to decompose a network’s parameter vector into a sum of other parameter vectors that correspond to distinct mechanisms of the network. The idea this definition expresses is that different mechanisms combine linearly in parameter space to form the whole network. If neural networks do consist of a unique set of mechanisms in a meaningful sense, the ability of APD to recover that set of mechanisms relies on the assumption that the mechanisms are encoded in the network parameters in this linear manner, at least up to some approximation. We call this the assumption of parameter linearity.
@@
Whether the assumption of parameter linearity is satisfied by all the non-toy neural networks that we might want to decompose is ultimately an empirical question. Current theoretical frameworks for computing arbitrary circuits in superposition (Vaintrob et al., 2024; Bushnaq and Mendel, 2024) do seem to satisfy this assumption. They linearly superpose mechanisms in parameter space to perform more computations than the model has neurons [^10] In such a manner that equation 9 should be satisfied up to terms scaling as ca. $\mathcal{O}(\epsilon)$, where $\epsilon$ is the noise level in the outputs of the target model due to superposition. This tentatively suggests that real models relying on superposition for computation do the same.
@@ A.2.1 Idealised loss: Minimum description length loss, $\mathcal{L}_{\text{MDL}}$
Optimizing our components $P_{c}$ to minimise $\mathcal{L}_{\text{MDL}}(x)$ would then yield a decomposition of the network that uses only small values for the total number of active components and the total rank of the active components on a particular forward pass.
@@ A.2.2 Practical loss: Top-$k$ formulation of APD
then $V_{c}$ and $U_{c}$ will be incentivised to effectively become proportional the right and left singular vectors for subnet $P_{c}$.
The Schatten norm of $P_{c}$ can then be written in factorised form as:
$\mathcal{L}_{\text{simplicity}}(x)$ $=\sum^{C}_{c=1}s_{c}(x)\sum_{l,m}\left(\sum_{i,j}U^{2}_{c,l,m,i}V^{2}_{c,l,m,j}\right)^{\frac{p_{2}}{2}}\,.$ (22)
@@ A.2.3 Alternative practical loss: APD formulation that uses an $L_{p}$ penalty on attributions
where $p_{1},p_{2}\leq 1.0$ are the norm of the attributions and the Schatten norm of the matrices respectively.

### v4

@@ 1 Introduction
To this end, we introduce Attribution-based Parameter Decomposition (APD), a method that decomposes neural network parameters into components that are optimized for these three properties. In brief, APD involves decomposing the parameter vector of any neural network into a sum of parameter components. They are optimized such that they sum to the target parameters while only a minimal number of them are necessary for the causal process that computes the network’s output for any given input. They are also optimized to be less complex individually than the entire network, in that they span as few directions in activation space as possible across all layers. APD can be understood as an instance of a broader class of Linear Parameter Decomposition (LPD) methods, a term we attempt to make precise in this paper.
@@ 2.2 Identifying networks’ mechanisms using Attribution-based Parameter Decomposition
APD aims to minimize the description length of the mechanistic components used by the network per data point over the training dataset. It decomposes the network’s parameters $\theta^{*}\in\mathbb{R}^{N}$ into a set of parameter components and directly optimizes them to be faithful, minimal, and simple. A discussion of how APD can be understood as an instance of a broader class of ‘linear parameter decomposition’ methods can be found in Appendix A.1, and more detailed discussion of how APD is based on the Minimum Description Length principle can be found in Appendix A.2.
@@
    Attribution step: We want to estimate the causal importance of each parameter component $P_{c}$ for the network’s output on each datapoint $f_{\theta^{*}}(x)$. In this step, we therefore calculate the attributions of each parameter component with respect to the outputs, $A_{c}(x)\in\mathbb{R}$. It would be infeasibly expensive to compute this exactly, since it would involve a large number of causal interventions, requiring one forward pass for every possible combination of component ablations. We therefore use an approximation. In this work, we use gradient attributions (Mozer and Smolensky, 1988; Molchanov et al., 2017; Nanda, 2022a; Syed et al., 2023), but other attribution methods may also work. This step therefore involves one forward pass with the target model to calculate the output and one backward pass per output dimension [^3] In the future, this could possibly be reduced to one backward pass with stochastic sources techniques (See Knechtli et al. (2016) (Chapter $3.6$) for an introduction to these methods). to compute the attributions with respect to the parameters, which are used to calculate the attributions with respect to each parameter component (Equation 10).
@@ 3.2 Toy Model of Compressed Computation
The target network is a residual MLP, consisting of a residual stream width of $d_{\rm resid}=1000$, a single MLP layer of width $d_{\rm mlp}=50$, a fixed, random embedding matrix with unit norm rows $W_{E}$, an unembedding matrix $W_{U}=W_{E}^{\top}$, and $100$ input features. See Figure 5 for an illustration of the network architecture. The large residual stream $d_{\rm resid}=1000$ was chosen as the trained target network performed better than the naive monosemantic baseline in this setting (small values of $d_{\rm resid}$ lead to higher interference and thus worse model performance). We chose fixed, instead of trained, embedding matrices to make it simpler to calculate the optimal monosemantic baseline and to simplify training.
@@ 4.1 Addressing issues that seem challenging from an activations-first perspective
    ‘Features as arbitrary functions’, but this fails to distinguish between features that appear to be fundamental abstractions (e.g. a ‘cat feature’) and those that don’t (e.g. a ‘cat OR car’ feature).
@@
In our work, we decompose neural networks into parameter components that minimize mechanistic description length, which we call the network’s ‘mechanisms’. Note that a network’s mechanisms are not equivalent to its ‘features’, but they might be related. Defining a network’s features as ‘properties of the input that activate particular mechanisms’ seems to overcome the definitional issues above.
In particular, it overcomes the issues in Definition 1 because a set of maximally faithful, minimal, and simple components should learn to correspond to ‘cat mechanisms’ and ‘car mechanisms’, but not ‘cat OR car mechanisms’ (unless the target network actually did have specific machinery dedicated to ‘cat OR car’ computations). APD also does not rely on a notion of human interpretability, overcoming the issue with Definition 2. It also seems to overcome the issues of Definition 3, since the definition is not circular and should also be able to identify multidimensional mechanisms (and hence multidimensional features that activate them), although we leave this for future work.
@@
Neural network representations appear not to neatly map to individual architectural components such as individual neurons, attention heads, or layers. Representations often appear to be spread across various architectural components, as in attention head superposition (Jermyn et al., 2023) or cross layer distributed representations (Yun et al., 2021; Lindsay et al., 2024). It is unclear how best to adapt SDL to each of these settings in order to tell concise mechanistic stories (Mathwin et al., 2024; Wynroe and Sharkey, 2024; Lindsay et al., 2024). A more general approach that requires no adaptation would be preferred. Interpretability in parameter space suggests a way to overcome this problem in general, since any architecture can in theory be decomposed into directions in parameter space without the method requiring adaptation.
@@ 5 Conclusion
This work introduces Attribution-based Parameter Decomposition (APD) as a fundamental shift in mechanistic interpretability: instead of analysing neural networks through their activations, we demonstrate that directly decomposing parameter space can reveal interpretable mechanisms that are faithful, minimal, and simple. Our approach suggests solutions to long-standing problems in mechanistic interpretability, including identifying minimal circuits in superposition, providing a conceptual foundation for ‘features’, enabling better understanding of ‘feature geometry’, and serving as an architecture-agnostic approach to neural network decomposition.
@@ Appendix A More detailed description of the APD method
A.1 Linear Parameter Decomposition
We want the components to be faithful to the original model in the sense that ideally, they should sum to form the target model’s parameters:
$\theta^{*}=\sum^{C}_{c=1}P_{c}\,.$ (6)
However, we also want most of the parameter components $P_{c}\in\mathbb{R}^{N}$ to not be ‘used’ on any one network input $x$, in the sense that we can ablate all but a few of them without changing the outputs of the network.
A definition of ‘inactive’ parameter components
We could try to operationalize the idea of most components being ‘inactive’ in the sense of playing no meaningful role in the computation of the output by initializing a new network with a parameter vector $\kappa(x)$ composed of only a few of the most ‘used’ parameter components, and demanding that:
@@ A.1 Linear Parameter Decomposition
A stricter definition of ‘inactive’ parameter components
If the ‘inactive’ components are not playing any meaningful role in the computation of the output, we should also be able to ablate or partially ablate them in any combination, and still get the same result. In general, we should get the same network output for any parameter configuration along any monotonic ‘ablation curve’ $\gamma_{c}(x,t)$ where $t\in[0,1]$:
$f(x|\kappa(x,t))\approx f(x|\theta^{*})\,.$
This is a stricter definition of ‘inactive’ that seeks to exclude cases like components $P_{c}$ and $P_{c^{\prime}}$ both being ‘active’ and important but canceling each other out. Without this stricter condition, we could have pathological solutions to the optimization problem. For example, suppose $P_{1},\dots,\,P_{C-1}$ are a large set of parameter vectors for expert networks specialized to particular tasks that a target network $\theta^{*}$ is capable of. However, also suppose that these parameter vectors are completely unrelated to the underlying mechanisms in the target network $\theta^{*}$. Then, we could set the last component to
$P_{C}=\theta^{*}-\sum^{C-1}_{c=1}P_{c}\,.$
Equations 6 and 9 together define what we mean when we say that we want to decompose a network’s parameter vector into a sum of other parameter vectors that correspond to distinct mechanisms of the network. The idea expressed in this definition is that different mechanisms combine linearly in parameter space to form the whole network. We call the class of methods that attempt to decompose neural network parameters into components that approximately satisfy Equations 6 and 9 Linear Parameter Decomposition (LPD) methods.
Component attributions
Directly checking that Equation 9 is satisfied would be computationally intractable. Instead, for APD, we try to estimate whether the condition is approximately satisfied by calculating attributions of the output to each component. Currently, we do this using gradient attributions (Finlayson et al., 2021; Molchanov et al., 2017; Nanda, 2022b). This estimates the effect of ablating $P_{c}$ as:
@@
If neural networks do consist of a unique set of mechanisms in a meaningful sense, the ability of APD and any other LPD methods to recover that set of mechanisms relies on the assumption that the mechanisms are encoded in the network parameters in the linear manner specified by equations 6 and 9, at least up to some approximation. We call this the assumption of parameter linearity.
@@
Whether the assumption of parameter linearity is satisfied by all the non-toy neural networks that we might want to decompose is ultimately an empirical question. Current theoretical frameworks for computing arbitrary circuits in superposition (Vaintrob et al., 2024; Bushnaq and Mendel, 2024) do seem to satisfy this assumption. They linearly superpose mechanisms in parameter space to perform more computations than the model has neurons [^11] In such a manner that equation 9 should be satisfied up to terms scaling as ca. $\mathcal{O}(\epsilon)$, where $\epsilon$ is the noise level in the outputs of the target model due to superposition. This provides a tentative theoretical basis to think that real models using computation in superposition do the same.
@@ A.2.1 Idealised loss: Minimum description length loss, $\mathcal{L}_{\text{MDL}}$
Optimizing our components $P_{c}$ to minimize $\mathcal{L}_{\text{MDL}}(x)$ would then yield a decomposition of the network that uses only small values for the total number of active components and the total rank of the active components on a particular forward pass.
@@ A.2.2 Practical loss: Top-$k$ formulation of APD
then $V_{c}$ and $U_{c}$ will be incentivised to effectively become proportional to the right and left singular vectors for subnet $P_{c}$.
The Schatten norm of $P_{c}$ can then be written in factorized form as:
$\mathcal{L}_{\text{simplicity}}(x)$ $=\sum^{C}_{c=1}s_{c}(x)\sum_{l,m}\left(\sum_{i,j}U^{2}_{c,l,m,i}V^{2}_{c,l,m,j}\right)^{\frac{p}{2}}\,.$ (22)
@@ A.2.3 Alternative practical loss: APD formulation that uses an $L_{p}$ penalty on attributions
where $p_{1},p_{2}\leq 1.0$ are the $p$-norms of the attributions and the Schatten norm of the matrices respectively.